OpenStudy (anonymous):

3 + a + b = a + 3 + b is an example of which algebraic property?

OpenStudy (zzr0ck3r):

its two, associative and commutative

OpenStudy (zzr0ck3r):

if it was (3 + a) + b = (a + 3) + b , then it would just be commutative

OpenStudy (zzr0ck3r):

but with no parenthesis, we need to use the fact that it does not matter what order we group our terms(this is called associative property)

OpenStudy (mathmate):

An example of associative would be: (a+b)+c = a+(b+c)

OpenStudy (zzr0ck3r):

But having said all that, I think the answer that they are looking for is commutative.

OpenStudy (zzr0ck3r):

@xspoon its commutative

OpenStudy (anonymous):

Thanks